OpenStudy (kt7779):

PLEASE HELP I NEED IT SOOOOOO BAD simplify 2m^2 - 2m + 3m^2

OpenStudy (anonymous):

5m^2-2m

OpenStudy (anonymous):

SOLUTION: m^3+2m^2-3m=0 Solve each equation so Far i have M(M^2+2m-3)=0

OpenStudy (anonymous):

are you asking for M?

OpenStudy (anonymous):

M=0, M=2/5

OpenStudy (anonymous):

thats wrong it is actually M=0, -3,1

OpenStudy (anonymous):

Do you need anymore help @kt7779

OpenStudy (kt7779):

i just need it simplified

OpenStudy (anonymous):

Simplified more? or just the 5m^2-2m

OpenStudy (kt7779):

i think that will work THANKS SO MUCH!

OpenStudy (anonymous):

welcome(: